A tennis ball player wins a match and hit the at 30m/s and at an angle of 40 above the horizontal on it way down was cough by a spectator 10m above the point was hit

1 . Calculator the time taken to reach the spectator

2. What is the magnitude of the velocity of the ball at impact (g .9.8m/s)


1
Expert's answer
2021-02-24T16:58:38-0500

1)

"h=v\sin{40}t-0.5gt^2\\10=30\sin{40}t-0.5(9.8)t^2\\t=0.61\ s"

2)


"v=\sqrt{(30\cos{40})^2+(30\sin{40}-9.8(0.61))^2}=26.6\frac{m}{s}"